OpenStudy (anonymous):

Applying the First Derivative Test (3/2)sin(πx/2-1) (a) find the critical numbers of f (b) find the open intervals on which the function is increasing or decreasing (c) apply the First Derivative Test to identify all relative extrema

OpenStudy (anonymous):

I was able to only get as far as the first derivative, but I get stuck trying to find the critical numbers

OpenStudy (anonymous):

Just to clear things up about what the function looks like: \[g(x)=\frac{ 3 }{ 2 }\sin(\frac{ \pi x }{ 2 }-1), [0,4]\]